OpenStudy (anonymous):

surface acrea of cyclinder Radius=10, height =21

OpenStudy (anonymous):

A≈1947.79

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

Finding the volume of a cylinder is the exact same formula! Step 1: Find the area of the base Step 2: Multiply it by the height In other words: Volume = (Area of base) • height Volume = (Area of a circle) • height

OpenStudy (anonymous):

area of base 3.14*100=314 ?

OpenStudy (ddcamp):

\[SA = 2\pi r^2 + 2 \pi r*h\]